### v1

Abstract
This paper develops a model-free framework for static fixed-income pricing and the replication of liability cash flows. We show that the absence of static arbitrage across a universe of fixed-income instruments is equivalent to the existence of a strictly positive discount curve that reproduces all observed market prices. We then study the replication and super-replication of liabilities and establish conditions ensuring the existence of least-cost super-replicating portfolios, including a rigorous interpretation of swap–repo replication within this static framework. The results provide a unified foundation for discount-curve construction and liability-driven investment, with direct relevance for economic capital assessment and regulatory practice.
Keywords: Fixed-Income Pricing; Static Arbitrage; Discount Curves; Liability Replication; Swap–Repo Replication
1 Introduction
The valuation and replication of insurance liabilities rely fundamentally on the structure of fixed-income markets. Regulatory frameworks such as Solvency II and the Swiss Solvency Test require insurers to discount expected liability cash flows using market-consistent yield curves and to assess hedging strategies based on portfolios of bonds, swaps, and repo transactions. [^3] For Solvency II see https://www.eiopa.europa.eu/browse/regulation-and-policy/solvency-ii_en. For the Swiss Solvency Test see https://www.finma.ch/en/supervision/insurers/cross-sectoral-tools/swiss-solvency-test-sst. Despite the operational importance of these practices, a basic structural question remains: under what conditions do the observed prices of fixed-income instruments define an arbitrage-free pricing system that admits a common discount curve and supports a consistent valuation and replication of expected liability cash flows?
This paper develops a general and model-free framework for fixed-income pricing based solely on static arbitrage. We show that the absence of (strict) arbitrage across a universe of fixed-income instruments is equivalent to the existence of a strictly positive (nonnegative) discount curve that reproduces all observed market prices. This result constitutes a fundamental theorem of fixed-income pricing and provides a theoretical foundation for discount-curve construction in insurance asset–liability management. To the best of our knowledge, this paper provides the first systematic treatment of static arbitrage in fixed-income price systems formulated directly in terms of dated cash flows and observed market prices, with an explicit focus on liability replication.
We then turn to the replication of expected insurance liability cash flows. Perfect replication typically fails when liability and asset cash-flow dates and cash flows do not align. This motivates a super-replication approach, for which we establish existence of a least-cost replicating portfolio and outline conditions ensuring well-posedness. A key insight is that swap–repo strategies naturally produce fixed-income cash-flow profiles, allowing swaps, repos, and floating-rate notes to be treated within the same static framework as coupon bonds.
The results presented here are consistent with standard fixed-income valuation and replication practices and clarify the conditions under which current regulatory discounting and liability valuation methodologies are well founded, as well as their structural limitations. They also point toward several avenues for further research, including uniqueness of super-replication portfolios, numerical methods for liability hedging, and implications for the construction of regulatory discount curves.
The theoretical foundations of arbitrage-free pricing are well established in the discrete-time and static asset pricing literature. Seminal contributions by [har_kre_79] and [dal_mor_wil_90] establish the equivalence between arbitrage-free price systems and the existence of linear pricing functionals, or state-price vectors; see [coc_01] and [foe_sch_16] for surveys and further references. These results apply to general financial markets and are not tailored to fixed-income price systems described directly in terms of dated cash flows. To the best of our knowledge, a systematic treatment of static arbitrage in fixed-income markets formulated explicitly at the level of cash-flow matrices and observed prices, and linked to the replication of expected liability cash flows, has not appeared in the literature.
The remainder of the paper is organized as follows. Section 2 develops the fundamental theorem of fixed-income pricing and establishes the equivalence between arbitrage-free prices and the existence of a discount curve. Section 3 applies this framework to the replication and super-replication of liabilities and derives conditions under which least-cost super-replication is feasible. Section 4 illustrates the theory through canonical fixed-income instruments, including coupon bonds, swaps, and repo-based strategies. Section 5 concludes with an outlook on open questions and directions for further research. All proofs are collected in the Appendix.
2 A Fundamental Theorem of Fixed-Income Pricing
We consider a universe of $M$ fixed-income instruments, indexed by $i=1,\dots,M$, with prices $P_{i}$ and deterministic cash flows $C_{i1},\dots,C_{iN}$ occurring at a common set of dates $0<x_{1}<\cdots<x_{N}$. [^4] If instrument $i$ has no cash flow at $x_{j}$, we set $C_{ij}=0$. We collect the prices in the $M\times 1$ vector $P$ and the cash flows in the $M\times N$ cash-flow matrix $C$.
We assume a frictionless market in which these instruments can be traded freely. A portfolio is represented by a vector $q\in\mathbb{R}^{M}$, where $q_{i}$ denotes the position in instrument $i$. Such a portfolio generates the cash-flow vector $q^{\top}C$ and has price $q^{\top}P$.
A central question, which has received surprisingly little attention in the literature, is whether the observed price system is internally consistent, in the sense that it does not permit static arbitrage across the fixed-income instruments. To address this question, we formalize the standard notions of static arbitrage in this setting. We begin with the most basic consistency requirement, namely that prices depend linearly on cash flows.
Theorem 2.1 (Law of one price).
The following statements are equivalent:
  1. (L1)
    The law of one price holds: every portfolio $q\in{\mathbb{R}}^{M}$ that generates zero cash flows, $q^{\top}C=0$, has zero price, $q^{\top}P=0$.
  2. (L2)
    There exists a real-valued discount curve $g:[0,\infty)\to{\mathbb{R}}$ with $g(0)=1$ and such that $P=C\,g(\bm{x})$.
This basic consistency property is strengthened by the following notions of arbitrage.
Definition 2.2.
A portfolio $q\in\mathbb{R}^{M}$ is called
  1. (i)
    a strict arbitrage if it has strictly negative price, $q^{\top}P<0$, and generates nonnegative cash flows, $q^{\top}C\geq 0$;
  2. (ii)
    an arbitrage if it has nonpositive price, $q^{\top}P\leq 0$, and generates nonnegative cash flows, $q^{\top}C\geq 0$, with at least one strictly positive component, that is, $(q^{\top}C)_{j}>0$ for some $j\leq N$.
We now derive the fundamental characterization of arbitrage in fixed-income pricing. To this end, define
$\mathcal{C}\coloneq\{\,C^{\top}q:q^{\top}P<0\,\},\qquad\overline{\mathcal{C}}\coloneq\{\,C^{\top}q:q^{\top}P\leq 0\,\},$
the sets of cash-flow vectors generated by portfolios with strictly negative and nonpositive prices, respectively. By construction, $\mathcal{C}$ is an open convex cone in $\mathbb{R}^{N}$, $\overline{\mathcal{C}}$ is a closed convex cone, and indeed $\overline{\mathcal{C}}=\operatorname{cl}(\mathcal{C})$.
The next two theorems characterize, first, the absence of strict arbitrage and, second, the absence of arbitrage, together with their implications for fixed-income pricing.
Theorem 2.3 (Absence of strict arbitrage).
The following statements are equivalent:
  1. (W1)
    There is no strict arbitrage.
  2. (W2)
    $\mathcal{C}\cap\mathbb{R}^{N}_{+}=\emptyset$.
  3. (W3)
    There exists a nonnegative discount curve $g\geq 0$ with $g(0)=1$ and such that $P=C\,g(\bm{x})$.
Moreover, each of the conditions (W1)(W3) implies the law of one price (L1)(L2) in Theorem 2.1.
Theorem 2.4 (Absence of arbitrage).
The following statements are equivalent:
  1. (S1)
    There is no arbitrage.
  2. (S2)
    $\overline{\mathcal{C}}\cap\mathbb{R}^{N}_{+}=\{0\}$.
  3. (S3)
    There exists a strictly positive discount curve $g>0$ with $g(0)=1$ and such that $P=C\,g(\bm{x})$.
Moreover, each of the conditions (S1)(S3) implies each of the weak conditions (W1)(W3) in Theorem 2.3.
Here is an example illustrating that no arbitrage is strictly stronger than no strict arbitrage.
Example 2.5.
Consider $M=N=2$ fixed-income instruments with prices $P_{1}=1$ and $P_{2}=c$, and cash-flow vectors $C_{1}=[1,0]$ and $C_{2}=[c,1]$, occurring at dates $0<x_{1}<x_{2}$, where $c>0$ is a fixed coupon rate. The pricing equations imply that any discount curve must satisfy $g(x_{1})=1$ and $g(x_{2})=0$. Thus the discount curve is nonnegative but not strictly positive. Consequently, there is no strict arbitrage, but arbitrage opportunities exist.
Indeed, the portfolio $q=[-c,\,1]^{\top}$ has price $q^{\top}P=0$ and generates the nonnegative, nonzero cash-flow vector $q^{\top}C=[0,1]$. Hence $q$ is an arbitrage but not a strict arbitrage.
3 Liability Replication and Super-Replication
We consider a liability portfolio that generates expected cash flows $Z=[Z_{1},\dots,Z_{N}]$ at the dates $x_{j}$. [^5] Expected cash flows are understood as expectations taken under the relevant risk-neutral forward measures. Under standard assumptions in insurance liability modeling, these expectations coincide with expectations under the real-world measure. [^6] As for the fixed-income instruments, $Z_{j}=0$ indicates that no expected liability cash flow occurs at date $x_{j}$. Our aim is to replicate these liability cash flows using a portfolio in the fixed-income instruments. In the strict sense, this means finding a portfolio $q\in\mathbb{R}^{M}$ such that
$q^{\top}C=Z.$ (1)
A portfolio $q$ satisfying (1) exists if and only if $Z^{\top}\in\operatorname{Im}(C^{\top})$, or equivalently, $\ker(C)\subseteq\ker(Z)$. In practice, this requirement may be too strong and need not hold. We therefore relax perfect replication to super-replication and define the super-replication set
$\mathcal{Q}\coloneqq\{\,q\in\mathbb{R}^{M}:q^{\top}C\geq Z\,\}.$
It is immediate that $\mathcal{Q}=\cap_{j\leq N}\{\,q\in\mathbb{R}^{M}:q^{\top}C_{:,j}\geq Z_{j}\,\}$ is a closed convex polyhedron in $\mathbb{R}^{M}$, and every portfolio satisfying (1) belongs to $\mathcal{Q}$.
The following result characterizes the feasibility of super-replication. It shows that the super-replication set $\mathcal{Q}$ is nonempty if and only if there exists no nonnegative discount vector that assigns zero value to all attainable cash flows while assigning strictly positive value to the liability cash flows $Z$.
Theorem 3.1 (Well-posedness).
The following statements are equivalent:
  1. (F1)
    Super-replication is feasible, that is, $\mathcal{Q}\neq\emptyset$.
  2. (F2)
    There exists no vector $v\in\mathbb{R}^{N}_{+}$ such that $Cv=0$ and $Zv>0$.
A sufficient condition for (F1)(F2) to hold is the existence of a fixed-income instrument $i$ with nonnegative cash flows whose effective cash-flow dates coincide with those of the liabilities, in the sense that $Z_{j}\neq 0$ if and only if $C_{ij}>0$ for all $j\leq N$.
In view of Theorem 3.1, super-replication is more likely to be feasible when the effective cash-flow dates of the fixed-income instruments and the liabilities are aligned. When such alignment fails, one may instead work with a proxy cash-flow matrix obtained by aggregating intermediate payments between liability dates. This proxy has a natural operational interpretation if intermediate cash proceeds can be held in a cash buffer, or, under no arbitrage, reinvested at forward-equivalent rates when the corresponding instruments are available.
Remark 3.2 (Cash-flow aggregation).
Let $x_{j_{1}}<\cdots<x_{j_{n}}$ denote the liability payment dates, that is, $Z_{j_{k}}\neq 0$. For each fixed-income instrument $i$, define a modified cash-flow matrix $\tilde{C}$ by aggregating payments between liability payment dates: with $j_{0}\coloneq 0$ and for $k=1,\dots,n$,
$\tilde{C}_{i,j}\coloneq\begin{cases}0,&j_{k-1}<j<j_{k},\\ \sum_{j_{k-1}<s\leq j_{k}}C_{i,s},&j=j_{k},\\ C_{i,j},&j>j_{n}.\end{cases}$
Operationally, this corresponds to assuming that intermediate cash flows can be held in a cash buffer until the next liability payment date.
Alternatively, under no arbitrage (S1), intermediate cash flows may be accumulated at forward-equivalent rates using a strictly positive discount curve $g>0$ by setting
$\tilde{C}_{i,j_{k}}=\sum_{j_{k-1}<s\leq j_{k}}\frac{g(x_{s})}{g(x_{j_{k}})}\,C_{i,s}.$
In this case $g$ is also a discount curve for the modified cash-flow matrix, $P=C\,g(\bm{x})=\tilde{C}\,g(\bm{x})$. This construction presumes the availability of the corresponding forward rate agreements.
We now seek a super-replication portfolio of minimal initial cost, that is, a solution to the linear program
$\min_{q\in\mathcal{Q}}q^{\top}P.$ (2)
The next theorem provides sufficient conditions under which a minimizer exists.
Theorem 3.3.
Assume that there is no arbitrage (S1), that $\ker(C^{\top})=\{0\}$, and that super-replication is feasible (F1). Then the optimization problem (2) admits at least one solution.
In summary, under the absence of arbitrage and mild regularity conditions, the super-replication problem is well posed: whenever super-replication is feasible, there exists a least-cost portfolio that dominates the liability cash flows.
Remark 3.4 (Alternative replication criteria).
An alternative to super-replication is quadratic hedging, which seeks a portfolio $q\in\mathbb{R}^{M}$ minimizing
$\|Z-q^{\top}C\|_{2}^{2}+\lambda\|q\|_{2}^{2},$
for some $\lambda>0$. This problem is strictly convex and therefore admits a unique solution. However, it penalizes over- and under-replication symmetrically, which is typically undesirable in liability replication. Asymmetric loss functions, such as smoothed hinge penalties, may provide a compromise between existence, uniqueness, and the one-sided nature of super-replication.
4 Illustrative Examples
We illustrate the framework with two prototypical classes of fixed-income instruments: coupon bonds and interest-rate swaps combined with repo financing.
4.1 Coupon Bonds
Government coupon bonds are the canonical fixed-income instruments in our setup. Such a bond is specified by a face value $F>0$ paid at maturity $T$ and coupon payments $c_{1},\dots,c_{n}$ at dates $0<T_{1}<\dots<T_{n}=T$. If the bond is indexed by $i$, its cash-flow vector is given by
$C_{i,j}=\begin{cases}c_{n}+F,&\text{if }x_{j}=T,\\ c_{k},&\text{if }x_{j}=T_{k},\ k=1,\dots,n-1,\\ 0,&\text{otherwise}.\end{cases}$ (3)
The corresponding price $P_{i}$ is observed in the bond market.
4.2 Interest-Rate Swaps and Repo
We next illustrate how fixed-income cash flows can be generated by combining interest-rate swaps with repo financing. [^7] Repo refers to a standard repurchase agreement.
Consider an overnight indexed receiver swap with fixed rate $R$, notional 1, and maturity $T=n\Delta$, for a fixed accrual period $\Delta>0$. [^8] For formal definitions of overnight indexed swaps (OIS), see, for example, [fil_tro_13, Section 2]. The holder receives fixed payments $\Delta R$ at dates $T_{k}=k\Delta\leq T$ and pays floating interest $\xi_{1},\dots,\xi_{m}$ at dates $0<t_{1}<\dots<t_{m}=T$, where $\xi_{j}$ is the repo rate earned over $t_{j-1},t_{j}]$.
We construct the following swap–repo strategy. Investing 1 in the repo market at inception and rolling it forward produces a unit payoff at $T$ and floating cash flows $\xi_{j}$ at the dates $t_{j}\leq T$. Entering the receiver swap at zero initial cost offsets these floating payments and yields fixed payments $\Delta R$ at $T_{k}\leq T$. Thus the combined strategy produces exactly the cash flows of a coupon bond with face value $F=1$ and coupons $c_{k}=\Delta R$. Under the absence of dynamic arbitrage, this synthetic bond must have initial price 1, since it is generated by a zero-cost swap combined with a unit repo investment. [^9] This notion of dynamic arbitrage is broader than the static definition in Definition 2.2. If the market price of the bond exceeded 1, one could short the bond and replicate it using the swap–repo strategy to lock in a risk-free profit, and similarly if the price were below 1. Hence the swap–repo strategy can be encoded in the form (3) with price $P_{i}=1$.
Remark 4.1 (Swap spread puzzle).
Under the absence of dynamic arbitrage, the swap–repo replication of coupon bonds implies that government bonds and interest rate swaps should be priced using the same discount curve. Empirically, however, government bond–swap spreads are persistently nonzero. This swap spread puzzle is commonly attributed to limits to arbitrage arising from market frictions, balance-sheet and regulatory constraints, and the safe-asset premium; see, for example, [chr_mir_21, wu_jar_24, aqu_etal_24].
To operationalize the swap–repo strategy in super-replication, suppose there are $M$ receiver swaps indexed by $i=1,\dots,M$. For simplicity we assume that the floating payment dates of the swaps and the repo roll dates coincide with the fixed-income grid, that is, $x_{j}=t_{j}$. Swap $i$ has maturity $n_{i}\Delta$ and fixed rate $R_{i}$.
We collect the fixed and floating swap payments in the $M\times N$ matrices $S$ and $\Xi$, defined by
$S_{i,j}=\begin{cases}\Delta R_{i},&\text{if }x_{j}=k\Delta,\ k=1,\dots,n_{i},\\ 0,&\text{otherwise},\end{cases}\qquad\Xi_{i,j}=\begin{cases}\xi_{j},&\text{if }x_{j}\leq n_{i}\Delta,\\ 0,&\text{otherwise}.\end{cases}$
We also define an $M\times N$ face-value matrix $F$ by
$F_{i,j}=\begin{cases}1,&\text{if }x_{j}=n_{i}\Delta,\\ 0,&\text{otherwise}.\end{cases}$
The coupon-bond cash-flow matrix corresponding to the swap–repo strategy is
$C=S+F=\underbrace{S-\Xi}_{\text{swap net cash flows}}+\underbrace{\Xi+F}_{\text{repo net cash flows}},$
and the initial prices of these synthetic bonds satisfy $P=\bm{1}$, since the swap has zero cost and the repo investment requires one unit of initial funding. [^10] The cash flows $\Xi+F$ may also be interpreted as those of floating-rate notes with matching payment and maturity dates.
Suppose that the assumptions of Theorem 3.3 are satisfied, and let $q$ be a minimizer of (2). The super-replication price of the liabilities is therefore $q^{\top}\bm{1}$. To implement the corresponding swap–repo strategy in practice, proceed as follows.
At inception $x_{0}=0$:
  • Invest (lend or borrow) the amount $q^{\top}\bm{1}$ in the repo market.
  • Enter $q_{i}$ units of receiver swap $i$ (zero cost).
At each payment date $x_{j}$:
  1. (i)
    Receive (or pay) floating interest and notional $\sum_{i:x_{j}\leq n_{i}\Delta}q_{i}(\xi_{j}+1)$ from the repo market.
  2. (ii)
    Reinvest $\sum_{i:x_{j}<n_{i}\Delta}q_{i}$ in the repo market.
  3. (iii)
    Offset the floating legs of the swaps, contributing $(q^{\top}\Xi)_{j}=\sum_{i:x_{j}\leq n_{i}\Delta}q_{i}\xi_{j}$.
  4. (iv)
    Receive (or pay) fixed swap interest $(q^{\top}S)_{j}$.
The repo-related transactions in (i)(ii) generate the cash flow $(q^{\top}(\Xi+F))_{j}$, while the swap-related transactions in (iii)(iv) generate $(q^{\top}(S-\Xi))_{j}$. Altogether, the net cash flow is $(q^{\top}(S+F))_{j}=(q^{\top}C)_{j}$, which by construction satisfies
$(q^{\top}C)_{j}\;\geq\;Z_{j},$
and thus super-replicates the expected liability cash flows.
5 Conclusion and Outlook
This paper develops a static, model-free framework for fixed-income pricing and the replication of liabilities. We establish a fundamental theorem of fixed-income pricing, characterizing arbitrage-free prices through the existence of a discount curve, and apply this framework to the super-replication of expected liability cash flows.
An important open question concerns the uniqueness of least-cost super-replicating portfolios. While existence follows under mild assumptions, the linear structure of the optimization problem generally permits multiple solutions. Identifying conditions under which uniqueness holds, or characterizing the structure of the optimal solution set, remains an interesting direction for further research.
Overall, the results provide a unified static foundation for fixed-income pricing and liability replication, with implications for asset–liability management, discount-curve construction, and economic and regulatory capital assessment.
Appendix A Proofs
This appendix collects all proofs.
A.1 Proof of Theorem 2.1
The law of one price is equivalent to $\ker(C^{\top})\subseteq\ker(P^{\top})$, which again is equivalent to $\operatorname{Im}P\in\operatorname{Im}C$.
[…]
A.2 Proof of Theorem 2.3
[…]
For the last statement, let $q$ be such that $q^{\top}C=0$. Now apply no strict arbitrage to $q$ and q-q, which implies $q^{\top}P\geq 0$ and $q^{\top}P\geq 0$, and thus $q^{\top}P=0$.
A.3 Proof of Theorem 2.4
[…]
The last statement follows trivially.
A.4 Proof of Theorem 3.1
[…]
The last statement follows by setting $q=te_{i}$, for $t=\frac{\max\{\max_{j}Z_{j},0\}}{\min_{j}C_{ij}}>0$ and where $e_{i}$ is the $i$th standard basis vector in ${\mathbb{R}}^{N}$.
A.5 Proof of Theorem 3.3
Let $q_{0}\in{\mathcal{Q}}$ and define the lower level set ${\mathcal{Q}}_{0}\coloneqq\{q\in{\mathcal{Q}}:q^{\top}P\leq q_{0}^{\top}P\}$ the set of all super-replicating strategies with price less than or equal to $q_{0}^{\top}P$. Then ${\mathcal{Q}}_{0}\neq\emptyset$ and we can replace $\mathcal{Q}$} by ${\mathcal{Q}}_{0}$ in (2). We now show that ${\mathcal{Q}}_{0}$ is compact, from which the theorem then follows.
Thereto, let $rec({\mathcal{Q}}_{0})=\{v\in{\mathbb{R}}^{M}:q+tv\in{\mathcal{Q}}_{0},\,\forall t\geq 0,\,\forall q\in{\mathcal{Q}}_{0}\}$ denote the recession cone of ${\mathcal{Q}}_{0}$. Let $v\in rec({\mathcal{Q}}_{0})$ and $q\in{\mathcal{Q}}_{0}$. Then it holds $(q+tv)^{\top}C\geq Z$ for all $t\geq 0$, which implies that $v^{\top}C\geq 0$. We claim that $v=0$. Indeed, suppose $v\neq 0$. Then $(v^{\top}C)_{j}>0$ for at least one cash flow $j\leq N$ by assumption. By no arbitrage therefore $v^{\top}P>0$ and thus $(q+tv)^{\top}P$ is unbounded from above in $t\geq 0$, which is absurd. Hence $rec({\mathcal{Q}}_{0})=\{0\}$.
On the other hand, ${\mathcal{Q}}_{0}$ is closed and convex by construction. Hence ${\mathcal{Q}}_{0}$ is compact, see [roc_97, Theorem 8.4], which completes the proof.